# Activated solids: Spontaneous deformations, non-affine fluctuations, softening, and failure

### v2

@@ IV.3 Hexatic order
This section examines the change in the hexatic order parameter to distinguish hexatic melting from solid melting. Each particle in the system can be associated with a hexatic bond orientational order. For the $k$-th particle, it is defined as $\psi_{6}^{k}=(1/n)\sum_{i=1}^{n}e^{i6\theta_{kj}}$, where $\theta_{kj}$ is the bond angle between particle $k$ and its topological (Voronoi) neighbor $j$ relative to the $x$-axis, and $n$ is the number of such neighbors. In Fig. 6(c), we plot the mean squared amplitude of the hexatic order parameter of the system
@@ IV.4 Topological defects
In Fig. 9(a),(b) we identify such defects in typical configurations across activated solid melting. We identify clusters of such defects, including five, seven, and eight fold defects. We use fraction of particles with $\nu$ neighbors, $n_{\nu}=\langle N_{\nu}\rangle/N$ to analyze the defect formation. Here $N_{\nu}$ indicates the number of particle with $\nu=5,6,7$ neighbors and $\langle\dots\rangle$ denotes the steady state average. Although larger departures from $\nu=6$ are present, their fraction is significantly low.
As activity $\Lambda$ exceeds $\Lambda=10.0$, the solid melting point, $\langle n_{6}\rangle$ decreases, while the defect fraction, $\langle n_{5,7}\rangle$, increases (see Fig. 9(c),(d)). The mean of total fraction of particles with non-six neighbors, $\delta_{d}=\frac{1}{N}\sum_{\nu\neq 6}N_{\nu}$, becomes non-zero at the solid melting point and increases with higher activity (Fig. 6(d)). The fluctuation in $\delta_{d}$ peaks twice: first at $\Lambda=10.2$, near the solid melting point, and again at $\Lambda=15.0$, at the hexatic melting point, reflecting enhanced defect fluctuations during both melting transitions (see inset of Fig. 6(d)).
@@ V Local softening: control of non-affinity and defect
The evolution of defect fraction in Fig. 10 (g-i) complements these observations. For $\Lambda=9$, defects intermittently form and annihilate; for $\Lambda\geq 10$, the defect fraction increases significantly within the activated region and more gradually outside it. The simultaneous rise of non-affinity and defect fraction suggests a coupled onset of mechanical disorder at these high activity.
@@ VI Discussion and Conclusion
In this work, we explored spontaneous deformations in active solids by quantifying local rearrangements using non-affine fluctuations relative to global strain. Our main achievements are three fold.
First, using scaling arguments and numerical simulations, we obtained precise predictions for the steady-state mean non-affinity as a function of activity parameters, such as active speed and persistence. The system-averaged non-affinity grows quadratically with active speed, scaling as $\Lambda^{2}$. When examined as a function of active persistence time, it shows an initial linear increase followed by saturation – behavior that cannot be captured by simple active-temperature estimates. Additionally, the non-affinity scales inversely with the distance from the melting density, $(\rho-\rho_{m})^{-1}$.
Second, at higher activity, the distribution of local non-affinity reveals the emergence of pronounced non-affine zones coexisting with regions of predominantly small non-affine fluctuations. As activity increases, the distributions broaden, become more skewed, and develop heavy tails, eventually exhibiting a secondary maximum indicative of coexisting low and high non-affinity regions. This coexistence precedes the divergence of non-affinity, which acts as a precursor to defect proliferation and eventual melting. Steady-state spatial correlations of local non-affinity exhibit a correlation length that grows with active speed and increases with persistence before saturating. The dynamical relaxation of system-averaged non-affinity is governed by the active persistence time.
Activity-induced softening is accompanied by a reduction in the shear modulus and the loss of both solid and hexatic order. The solid shows a two-step melting in a manner similar to the BKTHNY scenario, mediated by proliferation of topological defects.
@@
Third, we demonstrated that spatially patterned activation provides a direct route to locally induce non-affinity and mechanical softening. Unlike previous schemes requiring complex, configuration-dependent many-body forces, our approach is experimentally accessible: by selectively activating regions of an otherwise passive colloidal crystal – such as with optically controlled self-thermophoretic Janus particles – non-affine deformations can be reliably induced. This method directly tunes activity to control non-affinity, avoids real-time particle tracking, and leverages the system’s intrinsic active dynamics, offering a practical and experimentally feasible pathway for local, tunable mechanical control.
@@
More generally, anisotropic activation protocols may enable directional control of solid mechanics, opening avenues for spatially programmable mechanical responses. Beyond enriching our understanding of active solid mechanics, these insights contribute broadly to the design of adaptive metamaterials and may inform how biological systems dynamically regulate their rigidity through internally generated stresses.
@@ Appendix D Probability distributions of local non-affinity
Figure 13 compiles stacked probability distributions $P(\chi)$ of the local non-affinity $\chi$ on log-log axes over a broad range of active speed $\Lambda$ and rotational diffusivity of the heading direction $\tilde{D}_{r}$. Rows correspond to increasing activity, $\Lambda=1,4,5,6,7$, while columns span rotational diffusivities $\tilde{D}_{r}=0.010,0.025,0.100,0.250,1,3,10,30,100,500$. At fixed active speed (across each row), increasing $\tilde{D}_{r}$ has a nuanced impact on the distribution of non-affinity. At high enough active speed, e.g., $\Lambda=6$, extremely large persistence ($\tilde{D}_{r}=0.01$) keeps the distribution unimodal, whereas producing a tail towards higher $\chi$ over an intermediate range of persistence. While persistence helps to produce bimodality, too large a persistence can produce jamming, suppressing the large $\chi$ fluctuations. Beyond this regime, progressive decrease in persistence suppresses the large-$\chi$ tail, and the distributions take unimodal form at high $\tilde{D}_{r}$.
@@
Appendix E Obtaining shear modulus
Figure 14: (a) Shear stress, $\sigma_{xy}^{tot}$, vs. shear strain, $\mathcal{E}_{xy}=\gamma$, at $\Lambda=7$ for 24 ensembles without significant non-linearity or hysteresis. Different colors represent individual ensembles, with the black line showing the linear fit, where the slope gives the shear modulus, $\langle{\cal G}\rangle=105.7\pm 0.679$ in units of $\epsilon/\sigma^{2}$. (b) Linear response trajectories and (c) non-linear response trajectories exhibiting significant hysteresis and nonlinearity. Non-linear trajectories are excluded from shear modulus calculation. Raw data are shown as semi-transparent circular markers, with opaque square markers representing averages over 100 consecutive values of $\gamma$ and $\sigma^{\mathrm{tot}}_{xy}$. Markers are color-coded by shear direction: forward shear $0\rightarrow\gamma_{m}$ in red, backward shear $\gamma_{m}\rightarrow-\gamma_{m}$ in green, and $\gamma_{m}\rightarrow 0$ in blue. (d) Fraction of linear ensembles, $n_{\mathrm{lin}}/n_{\mathrm{tot}}$, as a function of activity $\Lambda$.
@@ Appendix E Obtaining shear modulus
We perform the calculations at $\tilde{D}_{t}=1.0$ and $\tilde{D}_{r}=3.0$. For $\Lambda\leq 6$, all four initial configurations exhibit linear stress-strain behavior. For $\Lambda>6$, non-linear deviations appear. We generate $n_{\mathrm{tot}}\approx 100$ initial configurations, discarding those with nonlinearity for shear modulus analysis, which becomes more common at higher $\Lambda$. The shear modulus is then computed by averaging the linear response data from the remaining $n_{\mathrm{lin}}$ hysteresis-free ensembles, as shown in Fig. 14(a). Examples of linear and non-linear ensembles are shown in Figs. 14(b) and (c).
The fraction $n_{\mathrm{lin}}/n_{\mathrm{tot}}$ decreases with $\Lambda$, reaching zero at the melting point, $\Lambda=10.0$ (Fig. 14(d)). This decline for $\Lambda>6.0$ indicates activity-driven softening, marked by a departure from linear elasticity. Near $\Lambda=10.0$, both $n_{\mathrm{lin}}$ and the shear modulus approach zero, signaling a fluidization transition with the complete loss of linear response.

### v3

@@ IV.3 Hexatic order
This section examines the change in the hexatic order parameter to distinguish hexatic melting from solid melting. Each particle in the system can be associated with a hexatic bond orientational order. For the $k$-th particle, it is defined as $\psi_{6}^{k}=(1/n)\sum_{i=1}^{n}e^{i6\theta_{kj}}$, where $\theta_{kj}$ is the bond angle between particle $k$ and its topological (Voronoi) neighbor $j$ relative to the $x$-axis, and $n$ is the number of such neighbors. In Fig. 6(c), we plot the mean squared amplitude of the hexatic order parameter of the system [34]
@@ IV.4 Topological defects
In Fig. 9(a),(b), we identify such defects in typical configurations across activated solid melting. We identify clusters of such defects, including five, seven, and eight-fold defects. We use the fraction of particles with $\nu$ neighbors, $n_{\nu}=\langle N_{\nu}\rangle/N$ to analyze the defect formation. Here $N_{\nu}$ indicates the number of particles with $\nu=5,6,7$ neighbors and $\langle\dots\rangle$ denotes the steady state average. Although larger departures from $\nu=6$ are present, their fraction is significantly low.
As activity $\Lambda$ exceeds $\Lambda=10.0$, the solid melting point, $\langle n_{6}\rangle$ decreases, while the defect fraction, $\langle n_{5,7}\rangle$, increases (see Fig. 9(c),(d)). The mean fraction of all particles with non-six neighbors, $\langle\delta_{d}\rangle=\frac{1}{N}\sum_{\nu\neq 6}\langle N_{\nu}\rangle$, becomes non-zero at the solid melting point and increases with higher activity (Fig. 6(d)). The fluctuation in $\delta_{d}$ peaks twice: first at $\Lambda=10.2$, near the solid melting point, and again at $\Lambda=15.0$, at the hexatic melting point, reflecting enhanced defect fluctuations at both the melting transitions (see inset of Fig. 6(d)).
@@ V Local softening: control of non-affinity and defect
The evolution of defect fraction in Fig. 10 (g-i) complements these observations. For $\Lambda=9$, defects intermittently form and annihilate; for $\Lambda\geq 10$, the defect fraction increases significantly within the activated region and more gradually outside it. The simultaneous rise of non-affinity and defect fraction suggests a coupled onset of mechanical disorder at these high activities.
@@ VI Discussion and Conclusion
In this work, we have investigated spontaneous deformations in active solids by quantifying local rearrangements through non-affine fluctuations relative to global strain. The observed scalings of the global non-affinity with activity parameters reflect the distinct mechanisms driving particle rearrangements in active solids. Increasing the self-propulsion speed enhances active fluctuations, thereby increasing non-affinity, while increasing persistence allows particles to maintain their heading over longer times, promoting more coherent local rearrangements. At low to moderate persistence, this leads to a rise in the mean global non-affinity, but at high persistence, the system becomes jammed, so further increases in persistence no longer enhance non-affinity, reflecting the interplay between persistent driving and structural constraints. These mechanisms differ fundamentally from equilibrium solids, where particle displacements undergo Gaussian fluctuations, resulting in chi-squared distributions of local non-affinity, with correlations decaying within roughly two lattice spacings. In contrast, the persistent local forces in active solids produce non-Gaussian fluctuations, longer-ranged non-affine correlations, and distributions of non-affinity that deviate markedly from the chi-squared form, revealing cooperative particle rearrangements absent in equilibrium.
Our main findings can be summarized in three key points. First, combining scaling arguments and numerical simulations, we quantified the steady-state mean non-affinity as a function of activity parameters. It grows quadratically with active speed and increases linearly with persistence before saturating, a behavior not captured by simple active-temperature estimates. Second, at higher activity, local non-affinity distribution broadens, becomes more skewed with heavy tails, and develops a secondary peak, reflecting the coexistence of rare large rearrangements and predominantly small fluctuations. This coexistence precedes the divergence of non-affinity, serving as a precursor to defect proliferation and eventual melting. Spatial correlations of non-affinity exhibit a correlation length that grows with active speed and increases with persistence before saturating, while the dynamical relaxation of the global non-affinity is set by the persistence time. Activity-induced softening reduces the shear modulus and destabilizes both solid and hexatic order, with two-step melting mediated by proliferation of topological defects. Non-affine fluctuations provide a sensitive probe of local rearrangements and defect precursors in crystalline solids, but their presence alone does not imply melting; rather, they signal regions where defect nucleation and subsequent structural transformations may occur.
@@
Third, we demonstrated that spatially patterned activation provides a direct and experimentally feasible route to locally induce non-affinity and mechanical softening. By selectively activating regions of a passive colloidal crystal – for example, via optically controlled self-thermophoretic Janus particles – non-affine deformations can be reliably induced without complex many-body interactions or real-time particle tracking. This approach allows tunable, localized control of mechanical response and opens the way for spatially programmable solid mechanics.
@@
Beyond deepening our understanding of active solid behavior, these results have broad implications for the design of adaptive metamaterials and may inform how biological systems dynamically regulate their rigidity through internally generated stresses.
@@ Appendix D Probability distributions of local non-affinity
Figure 13 compiles stacked probability distributions $P(\chi)$ of the local non-affinity $\chi$ on log-log axes over a broad range of active speed $\Lambda$ and rotational diffusivity of the heading direction $\tilde{D}_{r}$, for an athermal system ($\tilde{D}_{t}=0$). Rows correspond to increasing activity, $\Lambda=1,4,5,6,7$, while columns span rotational diffusivities $\tilde{D}_{r}=0.010,0.025,0.100,0.250,1,3,10,30,100,500$. At fixed active speed (across each row), increasing $\tilde{D}_{r}$ has a nuanced impact on the distribution of non-affinity. At high enough active speed, e.g., $\Lambda=6$, extremely large persistence ($\tilde{D}_{r}=0.01$) keeps the distribution unimodal, yet produces a tail towards higher $\chi$ over an intermediate range of persistence. While persistence helps produce bimodality, too large a persistence can lead to jamming, suppressing the large $\chi$ fluctuations. Beyond this regime, a progressive decrease in persistence suppresses the large-$\chi$ tail, and the distributions become unimodal at high $\tilde{D}_{r}$.
@@
Figure 14: (a) Representative time series of the global non-affinity $X$ at $\Lambda=6.0$ and $\tilde{D}_{r}=0.1$, showing the spontaneous emergence and decay of highly non-affine regions. (b) Configurations at the marked times (i)–(iii) in (a). Particles are colored by $\log_{10}(\chi_{i})$, with defects indicated as in Fig. 10(b,c). Panels (i) and (iii) show cropped (half-width) views around the event, while (ii) shows the full simulation box.
Figure 15: (a) Similar to Fig. 14, the time series of the global non-affinity $X$ at $\Lambda=10.0$ and $\tilde{D}_{r}=3.0$, configurations at the three marked times (i)–(iii) in the time series $X(t)$ shown at $t=87\,\tau_{u}$, $89.72\,\tau_{u}$, and $92\,\tau_{u}$, respectively. Particles are colored by $\log_{10}(\chi_{i})$, with defects indicated as in Fig. 10(b,c).
Appendix E Non-affine configurations
To better understand the bimodal distributions $P(\chi)$ in Fig. 3 and Fig. 13 for athermal systems ($\tilde{D}_{t}=0$), we show representative time series of $X(t)$ together with illustrative configurations for two parameter sets. The bimodality in $P(\chi)$ arises from both intermittent plastic bursts and steady-state coexistence of non-affine zones within an otherwise affine background. Fig. 14(a) shows a burst event at $\Lambda=6.0$ and $\tilde{D}_{r}=0.1$. The corresponding particle configurations in Fig. 14(b), colored by local non-affinity, resolve the sequence: nucleation of a localized high-$\chi$ region, growth into an extended slip-like band, and partial relaxation toward the low-$\chi$ background (Fig. 3(b)). In contrast, Fig. 15 illustrates the formation and growth of branched non-affine zones that eventually stabilize, producing steady-state coexistence of local non-affine rearrangements with an otherwise affine solid. These examples are illustrative: both transient bursts that fully relax and events that stabilize into persistent non-affine zones contribute to the bimodal distributions observed across the parameter regimes.
Appendix F Obtaining shear modulus
@@
We perform the calculations at $\tilde{D}_{t}=1.0$ and $\tilde{D}_{r}=3.0$. For $\Lambda\leq 6$, all four initial configurations exhibit linear stress-strain behavior. For $\Lambda>6$, non-linear deviations appear. We generate $n_{\mathrm{tot}}\approx 100$ initial configurations, discarding those with nonlinearity for shear modulus analysis, which becomes more common at higher $\Lambda$. The shear modulus is then computed by averaging the linear response data from the remaining $n_{\mathrm{lin}}$ hysteresis-free ensembles, as shown in Fig. 16(a). It displays the total shear stress, $\sigma_{xy}^{\mathrm{tot}}$, as a function of shear strain, $\mathcal{E}_{xy}=\gamma$, at $\Lambda=7$ for 24 independent ensembles. All trajectories exhibit a linear response without noticeable hysteresis. Different colors denote individual ensembles, while the black line is a linear fit whose slope yields the shear modulus, $\langle{\cal G}\rangle=105.7\pm 0.679$ (in units of $\epsilon/\sigma^{2}$). Here we reiterate that the mean swim stress remains unchanged under shear, as confirmed by numerical calculations; thus, the shear modulus is determined solely by the virial stress.
Figure 16: Shear-stress response under cyclic shear. (a) Total shear stress $\sigma_{xy}^{\mathrm{tot}}$ vs shear strain $\gamma$ at $\Lambda=7$ for 24 ensembles; the linear fit (black) yields $\langle{\cal G}\rangle=105.7\pm 0.679$ ($\epsilon/\sigma^{2}$). (b,c) Representative linear and nonlinear (hysteretic) trajectories; nonlinear cases are excluded from the modulus estimate. Symbols denote raw data (semi-transparent circles) and averages over 100 points (squares); colors indicate shear direction. (d) Fraction of linear ensembles, $n_{\mathrm{lin}}/n_{\mathrm{tot}}$, vs activity $\Lambda$.
Panels (b) and (c) in Fig. 16 show representative examples of linear-response trajectories and nonlinear trajectories exhibiting pronounced hysteresis and nonlinearity, respectively; the latter are excluded from the shear-modulus estimate. Raw data are shown as semi-transparent circles, while opaque squares represent averages over 100 consecutive values of $\gamma$ and $\sigma_{xy}^{\mathrm{tot}}$. Colors indicate the shear direction: forward shear $0\rightarrow\gamma_{m}$ (red), reverse shear $\gamma_{m}\rightarrow-\gamma_{m}$ (green), and return $\gamma_{m}\rightarrow 0$ (blue).
Fig. 16(d) shows the fraction of linear ensembles, $n_{\mathrm{lin}}/n_{\mathrm{tot}}$, as a function of the activity parameter $\Lambda$. It decreases with $\Lambda$, reaching zero at the melting point, $\Lambda=10.0$ (Fig. 16(d)). This decline for $\Lambda>6.0$ indicates activity-driven softening, marked by a departure from linear elasticity. Near $\Lambda=10.0$, both $n_{\mathrm{lin}}$ and the shear modulus approach zero, signaling a fluidization transition with the complete loss of linear response.